# Applied Antifragility in Natural Systems: Evolutionary Antifragility

### v1

@@ Abstract
This chapter introduces evolutionary antifragility as the time-scale interaction characteristics of a natural dynamic system. It describes the benefit derived from input distribution unevenness, based on the emergent system dynamics and its uncertain and volatile interactions with the operating environment described by unknown disturbances. We consider methods for the detection, analysis, and modelling of cancer, environmental, microbiota, and social systems antifragility. [^1] This is a preprint of the following work :::: Cristia Axenie, Roman Bauer, Oliver López Corona, Jeffrey West, Applied Antifragility in Natural Systems From Principles to Applications, Springer Nature reproduced with the permissions of Springer Nature. The final authenticated version is available online at https://link.springer.com/book/9783031903908

### v2

@@ Abstract
This chapter introduces evolutionary antifragility as the time-scale interaction characteristics of a natural dynamic system. It describes the benefit derived from input distribution unevenness, based on the emergent system dynamics and its uncertain and volatile interactions with the operating environment described by unknown disturbances. We consider methods for the detection, analysis, and modeling of cancer, environmental, microbiota, and social systems antifragility. [^1] This is a preprint of the following work :::: Cristian Axenie, Roman Bauer, Oliver López Corona, Jeffrey West, Applied Antifragility in Natural Systems From Principles to Applications, Springer Nature reproduced with the permissions of Springer Nature. The final authenticated version is available online at https://link.springer.com/book/9783031903908